OpenStudy (anonymous):

What is the slope of the line containing the points (2,6) and (-3,-4)? 2 -2 -10 -2/5

OpenStudy (anonymous):

m = (y' - y)/(x' - x) m = (6 - (-4)) / (2 - (-3)) m = 10/5 m = 2